OpenStudy (anonymous):

Solve D = 1 over 4(e – 3f) for e. You must show all work.

OpenStudy (anonymous):

D=1/4(e-3f) D=1/4e-12f 4e-12f=1/D 4e=1/D+12f e=1/4D+3f